The following data give the number of hours 5 students spent studying and their corresponding grades on their exams. Hours Studying 4 5 5 5 5 Grades 63 67 77 95 100 Step 1 of 5 : Calculate the sum of squared errors (SSE). Use the values b0=−24.0000 and b1=21.7500 for the calculations. Round your answer to three decimal places.
Answer #1

Solution

Back-up Theory

The linear regression model: Y = β0 + β1X + ε, ………………..………………………..(1)

Sum of squared errors (SSE) = Syy – b12.Sxx ..........................................................(2)

Where

Mean X = Xbar = (1/n) Σ(i = 1 to n)xi ………………………………………….……….….(3)

Mean Y = Ybar = (1/n) Σ(i = 1 to n)yi ………………………………………….……….….(4)

Sxx = Σ(i = 1 to n)(xi – Xbar)2 ………………………………………………..…………....(5)

Syy = Σ(i = 1 to n)(yi – Ybar)2 ……………………………………………..…….…………(6)

Now to work out the solution,

Let x = hours of study and y = grade

n

5

Xbar

4.80

ybar

80.40

Sxx

0.80

Syy

1091.20

b1

21.75

SSE

712.750

So, SSE = 712.750 Answer

DONE